# Streaming Max-Cut in General Metrics

### v1

@@ 4 An Algorithm for Sliding-Window Streams
The following reduction was established in [BO10], which reduces the task of devising sliding-window algorithms to insertion-only case, provided that the insertion-only algorithm is smooth.
@@
Our main technical lemma in this section is the following smoothness bound for $\operatorname{Max-Cut}$. Let $S$ be a subset of $P$. Here, we do not directly analyze the smoothness of $f(S)=\operatorname{Max-Cut}(S)$. Instead, we add a term $|S|\cdot\epsilon/n$ to the $\operatorname{Max-Cut}$ value, and this term is crucially needed. In fact, without this additional term, one can show that $f(S):=\operatorname{Max-Cut}(S)$ is not smooth. [^3] To see this, consider the 1D case, and $f$ does not satisfy Definition 4.2 for $A:=\{1,1\},B:=\{0\},C:=\{0\}$. The proof of Lemma 4.4 can be found in Section 4.1.
@@
Next, we show that the approximation ratio is $1+4\epsilon$. Specifically, let $W$ be the point set of an arbitrary sliding window, and we analyze the ratio for approximating $\operatorname{Max-Cut}(W)$. We start with the case $\operatorname{Max-Cut}(W)>0$, and we argue that in this case the algorithm must return $t$ as the return value instead of 0. To see this, since the minimum non-zero distance in the metric is 11, we get $\operatorname{Max-Cut}(W)\geq 1$ and hence $f(W)>\operatorname{Max-Cut}(W)\geq 1$. Since $t$ is a $(1+2\epsilon)$-approximation of $f(W)$, $t$ is still greater than $0.5$, and this implies that the return value of the algorithm is $t$. Now, to further bound the ratio of $t$, note that $|W|\cdot\epsilon/n\leq\epsilon$ and recall that we argued $\operatorname{Max-Cut}(W)\geq 1$, we know that $f(W)=\operatorname{Max-Cut}(W)+|W|\cdot\epsilon/n$ is $(1+\epsilon)$-approximation to $\operatorname{Max-Cut}(W)$. Since the return value $t$ is a $(1+2\epsilon)$-approximation to $f(S)$ by Lemma 4.3, we conclude that the return value $t$ is $(1+4\epsilon)$-approximation of $\operatorname{Max-Cut}(W)$ .
@@
Finally, the failure probability comes from Lemma 4.3, and it is at most $\frac{\epsilon}{192\log n}\cdot\frac{\log n}{\beta}\leq\frac{1}{3}$.
@@
Before we proceed to the proof, we need the following mathematical facts about the $\operatorname{Max-Cut}$ function.
@@
If $B=\emptyset$, then $f(A)=f(B)=0$, so $A=B=\emptyset$. We can directly get $f(A\cup C)=f(B\cup C)$. In the following, we assume that $B\neq\emptyset$.
To begin with, we assume that $\operatorname{cut}(B,B)>0$. By the premises of
$(1-\beta)f(A)\leq f(B),$
we have
@@
Note that $1-\beta\geq\frac{1}{2}$, we have
$\frac{1}{8}\operatorname{cut}(A\setminus B,A\setminus B)+\frac{1}{4}\operatorname{cut}(A\setminus B,B)\leq\beta f(B)$
As $|B|\cdot\epsilon/n\leq n\cdot\epsilon/n\leq 1\leq\operatorname{Max-Cut}(B)$, we get
@@
Summing all such $u\in A\setminus B$, we get
$\operatorname{cut}(A\setminus B,B)\geq\frac{|A\setminus B|}{2|B|}\operatorname{cut}(B,B)$
Combining this inequality with Equation 6, we have
@@
Next, fix some $c\in C$. By triangle inequality,
$0\leq\sum_{x\in A\setminus B}\sum_{y\in B}[\operatorname{dist}(x,y)+\operatorname{dist}(y,c)-\operatorname{dist}(x,c)]$
so
$|B|\cdot\sum_{x\in A\setminus B}\operatorname{dist}(x,c)\leq\operatorname{cut}(A\setminus B,B)+|A\setminus B|\sum_{y\in B}\operatorname{dist}(y,c)$
Combining this inequality with Equation 6, we get
@@
Combining (8) and (9),
$|B|\cdot\sum_{x\in A\setminus B}\operatorname{dist}(x,c)\leq 8\beta|B|\sum_{x\in B}\operatorname{dist}(x,c)+|A\setminus B|\sum_{x\in B}\operatorname{dist}(x,c)$
which implies
$\sum_{x\in A\setminus B}\operatorname{dist}(x,c)\leq 8\beta\sum_{x\in B}\operatorname{dist}(x,c)+\frac{|A\setminus B|}{|B|}\sum_{x\in B}\operatorname{dist}(x,c)\leq 16\beta\sum_{x\in B}\operatorname{dist}(x,c)$
where the second inequality is from (7). Summing over $c\in C$, we obtain
$\operatorname{cut}(A\setminus B,C)\leq 16\beta\operatorname{cut}(B,C)$
This together with (5) and (6) implies
@@
$\leq 12\beta\operatorname{cut}(B,B)+16\beta\operatorname{cut}(B,C)$
$\leq 16\beta\operatorname{cut}(B\cup C,B\cup C)$
$=\frac{1}{4}\alpha\operatorname{cut}(B\cup C,B\cup C)$
$\leq\alpha\operatorname{Max-Cut}(B\cup C),$
@@
Note that
$(1-\alpha)\cdot|A|\leq(1-8\beta)\cdot(1+8\beta)|B|\leq|B|$
Hence
$(1-\alpha)f(A\cup C)\leq f(B\cup C)$
This verifies Definition 4.2 and finishes the proof.
@@ 5 Lower Bound for Dynamic Streams
In this section, we establish the lower bound for the problem in the dynamic streaming setting. We begin by describing the dynamic streaming model.
Model
In the dynamic streaming mode, given an underlying metric space, a set of points $P\subseteq V$ in this metric is presented as a data stream of IDs insertions and deletes. At the end of the stream, the algorithm needs to report an estimation to the $\operatorname{Max-Cut}$ value for the set of points that are still not deleted. The algorithm has access to a distance oracle, which takes two IDs as input and returns the distance between the corresponding points. A query to the oracle is permitted only if both IDs are valid, meaning that they have appeared in the stream. If the algorithm queries the oracle with any invalid ID, it immediately fails.
In contrast to the model of [CSS16], which assumes that the smallest unit of operation is the point ID itself (and no encoding of IDs is allowed), here the algorithm may store or manipulate arbitrary encoding of IDs. Besides, [CSS16] requires that the algorithm can only query points that are currently stored by the algorithm, while we do not need this assumption.
@@
For any streaming algorithm $\mathcal{A}$ and sufficiently large integer $n$, there exists a metric space and dynamic point (ID) stream of length $n$ (in this metric) such that $\mathcal{A}$ must use $\Omega(n^{1/3})$ space to compute $\operatorname{poly}(n)$-approximation of $\operatorname{Max-Cut}$ with probability at least $0.55$.
@@
There exists an input distribution $\mathcal{D}$ over metric space and a dynamic point (ID) stream in this metric such that, for any deterministic algorithm $\mathcal{A}$, when its input is sampled from $\mathcal{D}$, $\mathcal{A}$ must use $\Omega(n^{1/3})$ space to compute $\operatorname{poly}(n)$-approximation of $\operatorname{Max-Cut}$ with probability at least $0.55$.
@@
For any randomized algorithm $\mathcal{A}$, let $\sigma$ denote the distribution of its random seed. Define $\mathcal{E}$ as the event that, on input $T$ sampled from $\mathcal{D}$ and seed $r\sim\sigma$, algorithm $\mathcal{A}$ outputs a $\operatorname{poly}(n)$-approximation using $o(n^{1/3})$ bits of space. We can write
@@
  • For each point $p_{i,j}$, define an identifier $\text{ID}_{i,j}:=(i,j,a_{i,j})$, where each $a_{i,j}$ is drawn independently and uniformly at random from $\{0,1\}$. Let $U:=\{(i,j,a_{i,j})\mid i\in[n^{\frac{2}{3}}],j\in[n^{\frac{1}{3}}]\}$ be the set of all IDs.
@@
Definition 5.3.
Given an input instance $T=(U,i^{*},j^{*},K)$ and an ID $(i,j,b)$, we say the ID is valid, if it belongs to $U$, i.e., $(i,j,b)\in U$; otherwise, it is invalid.
@@
The stream is constructed as follows:
@@
We will prove that no (deterministic) algorithm using less than $0.01\cdot n^{\frac{1}{3}}$ bits of space can compute an estimate $\eta>0$ that is, with high probability, a $\operatorname{poly}(n)$-approximation to the Max-Cut value of the remaining points. Let $m:=0.01\cdot n^{\frac{1}{3}}$.
@@
If $K=\Delta$, then the point $p_{i^{*},j^{*}}$ is an outlier: its distance to any other point is $\Delta$, while all other pairs are at distance 11. Thus, the maximum cut separates $p_{i^{*},j^{*}}$ from the rest, achieving cut value $\Delta\cdot(n^{\frac{1}{3}}-1)$. ∎
@@
By 5.4, it suffices to show that the algorithm cannot differentiate $K=1$ or $K=\Delta$. Recall that in (10), the algorithm learns $K$ only if the algorithm queries the distance between $p_{i^{*},j^{*}}$ and another point in $R_{i^{*}}$, since otherwise, the input to the algorithm would be the same, and the algorithm must give the same return value because the algorithm is deterministic, in which case the algorithm cannot differentiate the value of $K$. This is a random event because of the randomness of $i^{*}$ and $j^{*}$, and we define $\mathcal{E}$ as this crucial event, as follows.
@@
The main technical lemma is to show $\mathcal{E}$ happens with small probability, which is stated as follows whose proof is postponed to Section 5.1.1. and this would conclude Lemma 5.2.
@@
In conclusion, by Lemma 5.6 the algorithm cannot differentiate whether $K=1$ or $K=\Delta$ with probability $0.9$. By 5.4, which readily implies that the algorithm must suffer a $\operatorname{poly}(n)$ factor error, provided that $\Delta$ is set to high-degree polynomial of $n$. ∎
@@
Without loss of generality, we assume the algorithm always uses ID in $\{(i,j,\ell)\mid i\in[n^{\frac{2}{3}}],j\in[n^{\frac{1}{3}}],\ell\in\{0,1\}\}$, as otherwise, the algorithm queries an invalid ID and fails immediately.
@@
Observe that $\operatorname{dist}$ is random and it particularly depends on $i^{*}$ and $j^{*}$. Consequently, this potentially makes the internal state of the deterministic algorithm depending on $i^{*}$ and $j^{*}$, and this makes it difficult to analyze the event $\mathcal{E}$. Actually, this does not matter, we can avoid such dependence. Intuitively, consider the earliest time in the stream that $\mathcal{E}$ happens, then before this time step, the algorithm is oblivious to the value of $K$, and by (10), $\operatorname{dist}=K$ if and only if the distance between $p_{i^{*},j^{*}}$ and some other point in $R_{i^{*}}$ is queried, one can conclude that up to this point, the algorithm is also oblivious to $i^{*},j^{*}$.
@@
Now, we argue that $\Pr[X^{\prime}\vee\mathcal{Y}^{\prime}]=\Pr[\mathcal{X}\vee\mathcal{Y}]$, and this would conclude the lemma. To this end, we first argue that $X^{\prime}$ happens if and only if $X$ happens. A key observation is that $\operatorname{dist}$ and $\operatorname{dist}^{\prime}$ agrees on all point pairs that do not contain $p_{i^{*},j^{*}}$ (recalling (10) and (13)). Now, fix a realization of $i^{*}$ and $j^{*}$. as long as the algorithm does not query a pair that contains $p_{i^{*},j^{*}}$ (in the insertion stage), the respond of distance remain the same regardless of $\operatorname{dist}$ or $\operatorname{dist}^{\prime}$, and this further means the query sequence of the algorithm is the same since the algorithm is deterministic. This implies that the time step of the insertion stage such that $X$ or $X^{\prime}$ happen is the same, denoted as $T$. Therefore, if both $X$ and $X^{\prime}$ do not happen at $T$, then they must both happen immediately after $T$; moreover, if $X$ ($X^{\prime}$ resp.) happens on or before $T$, then $X^{\prime}$ ($X$ resp.) must also happen because the algorithm is deterministic and the input is the same. We conclude that $\Pr[X]=\mathbb{E}[\Pr[X\mid i^{*},j^{*}]]=\mathbb{E}[\Pr[X^{\prime}\mid i^{*},j^{*}]]=\Pr[X^{\prime}]$. Similarly, still condition on the realization of $i^{*},j^{*}$, if neither $X$ nor $X^{\prime}$ happens, then either $Y$ and $Y^{\prime}$ both happens or neither of the two happens. Hence, this concludes that $\Pr[\mathcal{X}^{\prime}\vee\mathcal{Y}^{\prime}]=\Pr[\mathcal{X}\vee\mathcal{Y}]$

### v2

@@ 4 Algorithms for Sliding-Window Streams
The following reduction was established in [BO10], which reduces sliding-window algorithm design to the insertion-only case, provided that the insertion-only algorithm is smooth.
@@
Our main technical lemma in this section is the following smoothness bound for $\operatorname{Max-Cut}$. Let $S$ be a subset of $P$. Here, we do not directly analyze the smoothness of $f(S)=\operatorname{Max-Cut}(S)$. Instead, we add a term $|S|\cdot\epsilon/n$ to the $\operatorname{Max-Cut}$ value, and this term is crucially needed. In fact, without this additional term, one can show that $f(S):=\operatorname{Max-Cut}(S)$ is not smooth (as we mention in Section 1.2). The proof of Lemma 4.4 can be found in Section 4.1.
@@
Next, we show that the approximation ratio is $1+4\epsilon$. Specifically, let $W$ be the point set of an arbitrary sliding window, and we analyze the ratio for approximating $\operatorname{Max-Cut}(W)$. We start with the case $\operatorname{Max-Cut}(W)>0$, and we argue that in this case the algorithm must return $t$ as the return value instead of 0. To see this, since the minimum non-zero distance in the metric is $1$, we get $\operatorname{Max-Cut}(W)\geq 1$ and hence $f(W)>\operatorname{Max-Cut}(W)\geq 1$. Since $t$ is a $(1+2\epsilon)$-approximation of $f(W)$, $t$ is still greater than $0.5$, and this implies that the return value of the algorithm is $t$. Now, to further bound the ratio of $t$, note that $|W|\cdot\epsilon/n\leq\epsilon$ and recall that we argued $\operatorname{Max-Cut}(W)\geq 1$, we know that $f(W)=\operatorname{Max-Cut}(W)+|W|\cdot\epsilon/n$ is $(1+\epsilon)$-approximation to $\operatorname{Max-Cut}(W)$. Since the return value $t$ is a $(1+2\epsilon)$-approximation to $f(S)$ by Lemma 4.3, we conclude that the return value $t$ is $(1+4\epsilon)$-approximation of $\operatorname{Max-Cut}(W)$.
@@
By Lemma 4.3, the failure probability is at most $\frac{\epsilon}{192\log n}\cdot\frac{\log n}{\beta}\leq\frac{1}{3}$. ∎
@@
Before we proceed, we need the following mathematical facts about $\operatorname{Max-Cut}$ function.
@@
If $B=\emptyset$, then $f(A)=f(B)=0$, so $A=B=\emptyset$. We can directly get $f(A\cup C)=f(B\cup C)$. In the following, we assume that $B\neq\emptyset$. To begin with, we assume that $\operatorname{cut}(B,B)>0$. By the premises of $(1-\beta)f(A)\leq f(B)$, we have
@@
Note that $1-\beta\geq\frac{1}{2}$, we have $\frac{1}{8}\operatorname{cut}(A\setminus B,A\setminus B)+\frac{1}{4}\operatorname{cut}(A\setminus B,B)\leq\beta f(B)$. As $|B|\cdot\epsilon/n\leq n\cdot\epsilon/n\leq 1\leq\operatorname{Max-Cut}(B)$, we get
@@
Summing all such $u\in A\setminus B$, we get $\operatorname{cut}(A\setminus B,B)\geq\frac{|A\setminus B|}{2|B|}\operatorname{cut}(B,B)$. Combining this inequality with Equation 12, we have
@@
Next, fix some $c\in C$. By triangle inequality, $0\leq\sum_{x\in A\setminus B}\sum_{y\in B}[\operatorname{dist}(x,y)+\operatorname{dist}(y,c)-\operatorname{dist}(x,c)]$ so $|B|\cdot\sum_{x\in A\setminus B}\operatorname{dist}(x,c)\leq\operatorname{cut}(A\setminus B,B)+|A\setminus B|\sum_{y\in B}\operatorname{dist}(y,c).$ Combining this inequality with Equation 12, we get
@@
Combining (14) and (15), it holds that
$|B|\cdot\sum_{x\in A\setminus B}\operatorname{dist}(x,c)\leq 8\beta|B|\sum_{x\in B}\operatorname{dist}(x,c)+|A\setminus B|\sum_{x\in B}\operatorname{dist}(x,c),$
which implies that
$\sum_{x\in A\setminus B}\operatorname{dist}(x,c)\leq 8\beta\sum_{x\in B}\operatorname{dist}(x,c)+\frac{|A\setminus B|}{|B|}\sum_{x\in B}\operatorname{dist}(x,c)\leq 16\beta\sum_{x\in B}\operatorname{dist}(x,c),$
where the second inequality is from (13). Summing over $c\in C$, we obtain $\operatorname{cut}(A\setminus B,C)\leq 16\beta\operatorname{cut}(B,C).$ This together with (11) and (12) implies
@@
[…]
$=\frac{1}{4}\alpha\operatorname{cut}(B\cup C,B\cup C)\leq\alpha\operatorname{Max-Cut}(B\cup C),$
@@
Note that $(1-\alpha)\cdot|A|\leq(1-8\beta)\cdot(1+8\beta)|B|\leq|B|$. Hence, $(1-\alpha)f(A\cup C)\leq f(B\cup C).$ This verifies Definition 4.2 and finishes the proof of Lemma 4.4.∎
@@ 5 Lower Bound for Dynamic Streams
In this section, we establish the lower bound for the problem in the dynamic streaming setting. We use the same model as in our upper bounds, specifically the one defined in Definition 1.1.
@@
For any streaming algorithm $\mathcal{A}$ and sufficiently large integer $n$, there exists a metric space and dynamic point (ID) stream, both of size $O(n)$, such that $\mathcal{A}$ must use $\Omega(n^{1/3})$ space to compute $O(\Delta/n^{1/3})$-approximation of $\operatorname{Max-Cut}$ with probability at least $0.55$ where $\Delta$ is the aspect ratio of the dataset.
@@
There exists an input distribution $\mathcal{D}$, over the metric space and a dynamic point (ID) stream in this metric, both of size $O(n)$, such that for any deterministic algorithm $\mathcal{A}$, when its input is sampled from $\mathcal{D}$, $\mathcal{A}$ must use $\Omega(n^{1/3})$ space to compute $O(\Delta/n^{1/3})$-approximation of $\operatorname{Max-Cut}$ with probability at least $0.55$.
@@
For any randomized algorithm $\mathcal{A}$, let $\sigma$ denote the distribution of its random seed. Define $\mathcal{E}$ as the event that, on input $T$ sampled from $\mathcal{D}$ and seed $r\sim\sigma$, algorithm $\mathcal{A}$ outputs an $O(\Delta/n^{1/3})$-approximation using $o(n^{1/3})$ bits of space. We can write
@@
  • For each point $p_{i,j}$, define an identifier $\text{ID}_{i,j}:=(i,j,a_{i,j})$, where each $a_{i,j}$ is drawn independently and uniformly at random from $\{0,1\}$. Let $U:=\{(i,j,a_{i,j})\mid i\in[n^{\frac{2}{3}}],j\in[n^{\frac{1}{3}}]\}$ be the set of all IDs. Note that there is a one-to-one correspondence between $U$ and $P$.
@@
We describe the stream as follows, where we use the language of points in $P$ but in fact they are presented to the algorithm as the corresponding point in $U$ (recalling that there is a one-to-one correspondence between $U$ and $P$):
@@
We will prove that no (deterministic) algorithm using less than $0.01\cdot n^{\frac{1}{3}}$ bits of space can compute an estimate $\eta>0$ that is, with high probability, an $O(\Delta/n^{1/3})$-approximation to the Max-Cut value of the remaining points. Let $m:=0.01\cdot n^{\frac{1}{3}}$.
@@
If $K=\Delta$, then the point $p_{i^{*},j^{*}}$ is an outlier: its distance to any other point is $\Delta$, while all other pairs are at distance $1$. Thus, the maximum cut separates $p_{i^{*},j^{*}}$ from the rest, achieving cut value $\Delta\cdot(n^{\frac{1}{3}}-1)$. ∎
@@
By Fact 5.3, it suffices to show that the algorithm cannot differentiate $K=1$ or $K=\Delta$. Recall that in (16), the algorithm learns $K$ only if the algorithm queries the distance between $p_{i^{*},j^{*}}$ and another point in $R_{i^{*}}$, since otherwise, the input to the algorithm would be the same, and the algorithm must give the same return value because the algorithm is deterministic, in which case the algorithm cannot differentiate the value of $K$. This is a random event because of the randomness of $i^{*}$ and $j^{*}$, and we define $\mathcal{E}$ as this crucial event, as follows.
@@
The main technical lemma is to show $\mathcal{E}$ happens with small probability, which is stated as follows whose proof is postponed to Section 5.1.1, and this would conclude Lemma 5.2.
@@
In conclusion, by Lemma 5.5 the algorithm cannot differentiate whether $K=1$ or $K=\Delta$ with probability $0.9$. By Fact 5.3, which readily implies that the algorithm must suffer a $\operatorname{poly}(n)$ factor error, provided that $\Delta$ is set to high-degree polynomial of $n$. ∎
@@
Observe that $\operatorname{dist}$ is random and it particularly depends on $i^{*}$ and $j^{*}$. Consequently, this potentially makes the internal state of the deterministic algorithm depending on $i^{*}$ and $j^{*}$, and this makes it difficult to analyze the event $\mathcal{E}$. We will see that this does not matter and there is a clean way around it. Intuitively, consider the earliest time in the stream that $\mathcal{E}$ happens, then before this time step, the algorithm is oblivious to the value of $K$, and by (16), $\operatorname{dist}=K$ if and only if the distance between $p_{i^{*},j^{*}}$ and some other point in $R_{i^{*}}$ is queried, one can conclude that up to this point, the algorithm is also oblivious to $i^{*},j^{*}$.
@@
Now, we argue that $\Pr[X^{\prime}\vee\mathcal{Y}^{\prime}]=\Pr[\mathcal{X}\vee\mathcal{Y}]$, and this would conclude the lemma. To this end, we first argue that $X^{\prime}$ happens if and only if $X$ happens. A key observation is that $\operatorname{dist}$ and $\operatorname{dist}^{\prime}$ agrees on all point pairs that do not contain $p_{i^{*},j^{*}}$ (recalling (16) and (19)). Now, fix a realization of $i^{*}$ and $j^{*}$. As long as the algorithm does not query a pair that contains $p_{i^{*},j^{*}}$ (in the insertion stage), the response of distance remain the same regardless of $\operatorname{dist}$ or $\operatorname{dist}^{\prime}$, and this further means the query sequence of the algorithm is the same since the algorithm is deterministic. This implies that the time step of the insertion stage such that $X$ or $X^{\prime}$ happen is the same, denoted as $T$. Therefore, if both $X$ and $X^{\prime}$ do not happen at $T$, then they must both happen immediately after $T$; moreover, if $X$ ($X^{\prime}$ resp.) happens on or before $T$, then $X^{\prime}$ ($X$ resp.) must also happen because the algorithm is deterministic and the input is the same. We conclude that $\Pr[X]=\mathbb{E}[\Pr[X\mid i^{*},j^{*}]]=\mathbb{E}[\Pr[X^{\prime}\mid i^{*},j^{*}]]=\Pr[X^{\prime}]$. Similarly, still condition on the realization of $i^{*},j^{*}$, if neither $X$ nor $X^{\prime}$ happens, then either $Y$ and $Y^{\prime}$ both happens or neither of the two happens. Hence, this concludes that $\Pr[\mathcal{X}^{\prime}\vee\mathcal{Y}^{\prime}]=\Pr[\mathcal{X}\vee\mathcal{Y}]$